Misty has $2.35 in dimes and nickels. She has one less nickel than dimes. How many of each coin does she have?

Answer by checkley71(8403)

.10X+.05(X-1)=2.35
.10X+.05X-.05=2.35
.15X=2.35+.05
.15X=2.40
X=2.40/.15
X=16 FOR THE NUMBER OF DIMES.
16-1=15 FOR THE NUMBER OF NICKELS.
PROOF
.10*16+.05*15=2.35
1.60+.75=2.35
2.35=2.35